Western Carolina University Dept. of Chemistry & Physics PHYS 230-01: General Physies I Instructor: Dr. Mesfin Woldeyohannes Quiz # 2 (September 08, 2017) Time Allowed: 15 Minutes NAME: 1) Uniformly Accelerated Motion I (10 Points) A car travelling at 25 mi/h is to stop on a 35-m long shoulder of the road (a) What is the required magnitude of the minimum deceleration? b)How much time will elapse during this minimum deceleration until the car stops? I want to answer question 1 a and b
Answer #1

Initial velocity, u= 25 mi/h= 25 \times \frac{1609.34m}{60 \times 60 s}= 11.176 m/s

Length of road covered, D= 35m

Final velocity, v= 0m/s (Since finally the car stops)

a). Let deceleration be "a", then using Equations of motion,

v^2-u^2= 2aD

(0)^2-(11.176)^2= 2a(35)

  a= -1.78m/s^2(ANS) (Negative sign shows deceleration)

b). Let time taken be "t", then using equations of motion

v= u+at

0= 11.176+(-1.78)t

  t= 6.278s(ANS)